# What Matters More For In-Context Learning under Matched Compute Budgets: Pretraining on Natural Text or Incorporating Targeted Synthetic Examples?

### v1

@@ 6 Conclusion and Future Work
We posed a single matched-compute question: At iso-FLOPs, does pure natural-text pretraining outperform a curriculum that explicitly targets the induction circuit via synthetic copy snippets? Using Bi-Induct (with forward, backward, or balanced directional copy injections), we evaluated models from 0.13B to 1B parameters, combining few-shot ICL performance with head-level telemetry and held-out perplexity as diagnostic tools.
Key finding: Natural data yields load-bearing heads, Bi-Induct induces more distributed, redundant heads. Bi-Induct consistently accelerates and broadens induction-head activity. In contrast, the natural-only baseline relies on a smaller subset of highly active heads. As can be seen from Figure 3 and Table 3, ablating top induction heads leads to larger performance drops in the natural-only model (evidence that its induction heads are more load-bearing), whereas Bi-Induct spreads responsibility across more redundant ‘backup’ heads.
Implications and Future Directions: Acceleration is not the whole story: what matters is whether the mechanism becomes load-bearing for task success. Our results align with previous reports that induction heads can evolve into ‘function heads’ that are causally necessary for ICL performance (Yin & Steinhardt, 2025). This helps explain why the baseline—despite slower emergence in most of our runs, and a lack of redundancy in all—can achieve stronger endpoints: its induction circuitry is more load-bearing.
The above points to a shift in focus: from inducing activity earlier to ensuring that such activity is necessary for the task. To support this, rather than relying solely on across-scale comparisons, future work should prioritize within-scale trajectory analysis to identify when behaviors become load-bearing, crucial for connecting data rewrites to mechanistic insights. While one might speculate that a stronger Bi-Induct schedule could push induction behavior towards more load-bearing circuits, that was not our design goal. We intentionally used a light, annealed curriculum to preserve LM usability (competitive perplexity) while probing ICL endpoints. Substantially heavier synthetic schedules could plausibly succeed in forcing functional induction, but at the cost of degraded language modeling quality and less plausible training data mixes. Accordingly, our claims are scoped to realistic, usability-oriented pretraining where synthetic data is a small perturbation annealed early. We would argue that future work should be similarly scoped.
@@ B.2 Mechanistic telemetry
We measure two equality-based copy circuits—induction and anti-induction— highlighted in prior work(e.g., (Olsson et al., 2022; Veitsman et al., 2025)).666See also (Wang et al., 2025; Yin & Steinhardt, 2025).. In a left-to-right causal decoder, attention flows from the later span back to the earlier span. Consider a repeated sequence with :
@@ C.1 Span length
We begin by testing how the snippet span affects outcomes. At 0.13B, with a fixed initial mix of linearly annealed over the full token budget, we sweep and report two endpoints of practical interest—(i) a 5-shot ICL composite over five standard LM benchmarks and (ii) held-out LM perplexity (PPL). We defer the function-probe suite of Todd et al. (2024) to the cross-scale experiments, where relative differences are more interpretable and the added compute is justified; the span-length results for this subsection are summarized in Table 6.
Across curricula, is a stable operating point that balances ICL and calibration: for Induction, ICL / PPL (vs. at and at ); for Anti, (vs. at , at ) respectively; for Balanced, (vs. at , at ). Very short spans () underperform on the ICL composite, while very long spans () offer no consistent ICL gain and tend to slightly worsen PPL. Hence we adopt for the remaining experiments. Beyond the ICL / PPL balance, shorter spans are operationally attractive: they minimize snippet length , which reduces potential overhead and makes it easier to pack snippets alongside natural sequences to exploit variable-length kernels—yielding compute savings when such packing is enabled.888All reported results are at iso-FLOPs; we do not take packing credits in our comparisons. Packing is a deployment optimization, not part of the evaluation protocol.
@@ C.2 Mix Ratio
Takeaways: Baseline naturally forms a few strong induction heads. Adding snippets increases the number of specialized induction heads up to moderate mixes (); higher mixes reduce spread and push peaks deeper. None of the curricula—especially Anti—produces meaningful anti-induction heads.
@@ D.2.2 Function-Probe Task Stress Tests
Because stress tests were explicitly considered during the development of the Todd et al. (2024) function-probe suite, we evaluate two that directly target ICL robustness: (i) label permutation within the in-context demonstration shots, and (ii) decision-rule sensitivity, contrasting the commonly reported HITS@3 with our primary metric, HITS@1. These tests probe robustness to spurious label–token mappings and to the choice of evaluation rule, respectively.
@@
Why permutation hurts 0.5B-1B but not 0.13B? At 0.13B, the robustness to label permutation suggests it benefits from demonstrations via format/topical priming and answer-frequency priors, but does not reliably exploit the labeltoken mapping. In contrast, 0.5B-1B models more strongly use the in-context mapping; permuting labels therefore contradicts a cue they have learned to trust, producing clear drops. This is consistent with reports that (i) labels in demos can be non-essential for smaller/less capable settings—format and priors often dominate (Min et al., 2022; Zhao et al., 2021), and (ii) the ability to override priors and follow contradictory, flipped labels emerges with scale (Wei et al., 2023).
@@
Following common practice for function-probe tasks in the Todd et al. (2024) suite—which reports HITS@3 (top-3 token accuracy)—we compare our primary decision rule (HITS@1) with HITS@3. ICL performance is summarized in Figure 7, and per-task HITS@3 accuracies are listed in Table 11. While HITS@3 increases absolute scores across the board, the relative ordering and gaps between variants remain effectively unchanged.

### v2

@@ 6 Conclusion and Future Work
We asked a single matched-compute question: for in-context learning, is it more effective to pretrain purely on natural text, or to allocate a small early-training budget to synthetic directional copy snippets that explicitly exercise the induction circuit? Using Bi-Induct (forward, backward, or balanced injections), we evaluated 0.13B–1B models with three complementary readouts: few-shot ICL performance, head-level copy telemetry, and held-out perplexity.
Bi-Induct reliably amplifies induction signatures, but it does not consistently improve few-shot ICL, and at 1B the natural-only baseline remains more load-bearing. For data-centric foundation model design, the broader lesson is methodological. Synthetic data interventions should not be evaluated only by whether they amplify a desired mechanistic signature. They should also be tested for whether that mechanism becomes causally necessary for the downstream behaviors of interest, and whether the intervention preserves natural-language modeling quality. In our study, Bi-Induct is best understood not as a generally superior curriculum, but as a controlled example showing that signature amplification alone is too weak a success criterion.
Several directions could extend this study. First, richer synthetic signals that incorporate semantic or linguistic structure may better align with the mechanisms underlying real-world induction behavior than the minimal token-level snippets used here. Second, scaling the analysis to larger models and longer context windows would clarify whether the emergence–vs.–load-bearing distinction persists in regimes where long-context retrieval becomes more critical. Finally, extending this methodology to alternative architectures, such as linear-attention models, may provide insight into whether induction-like behaviors in those systems become structurally necessary for long-context reasoning or remain redundant artifacts of training.
Limitations
Our conclusions are specific to a lightweight copy-based intervention under the conditions studied here. First, the injected snippets are token-level and intentionally minimal; they are not tied to richer linguistic or semantic structure, so our findings should not be read as ruling out more expressive mechanism-aware data rewrites. Second, our study uses final-checkpoint mechanistic analysis and models up to 1B parameters; broader generalization to larger scales remains open. Third, our runs use a context length of 1,024, so we do not claim that the same trade-offs must hold in substantially longer-context settings, where induction-like retrieval may matter more. Finally, because Bi-Induct replaces a fraction of natural data under iso-FLOPs, some of the observed trade-off may reflect natural-text displacement in addition to mechanistic redundancy.
These limitations sharpen, rather than weaken, the main takeaway: mechanism-aware data design should be assessed against a stricter standard than whether it makes a target circuit easier to measure. The key question is whether it makes that circuit matter.
@@ B.2 Mechanistic telemetry
We measure two equality-based copy circuits-induction and anti-induction- highlighted in prior work(e.g., (Olsson et al., 2022; Veitsman et al., 2025)).666See also (Wang et al., 2025; Yin and Steinhardt, 2025).. In a left-to-right causal decoder, attention flows from the later span back to the earlier span. Consider a repeated sequence with :
@@ C.1 Span length
We begin by testing how the snippet span affects outcomes. At 0.13B, with a fixed initial mix of linearly annealed over the full token budget, we sweep and report two endpoints of practical interest-(i) a 5-shot ICL composite over five standard LM benchmarks and (ii) held-out LM perplexity (PPL). We defer the function-probe suite of Todd et al. (2024) to the cross-scale experiments, where relative differences are more interpretable and the added compute is justified; the span-length results for this subsection are summarized in Table 6.
Across curricula, is a stable operating point that balances ICL and calibration: for Induction, ICL / PPL (vs. at and at ); for Anti, (vs. at , at ) respectively; for Balanced, (vs. at , at ). Very short spans () underperform on the ICL composite, while very long spans () offer no consistent ICL gain and tend to slightly worsen PPL. Hence we adopt for the remaining experiments. Beyond the ICL / PPL balance, shorter spans are operationally attractive: they minimize snippet length , which reduces potential overhead and makes it easier to pack snippets alongside natural sequences to exploit variable-length kernels-yielding compute savings when such packing is enabled.888All reported results are at iso-FLOPs; we do not take packing credits in our comparisons. Packing is a deployment optimization, not part of the evaluation protocol.
@@ C.2 Mix Ratio
Takeaways: Baseline naturally forms a few strong induction heads. Adding snippets increases the number of specialized induction heads up to moderate mixes (); higher mixes reduce spread and push peaks deeper. None of the curricula-especially Anti-produces meaningful anti-induction heads.
@@ D.2.2 Function-Probe Task Stress Tests
Because stress tests were explicitly considered during the development of the Todd et al. (2024) function-probe suite, we evaluate two that directly target ICL robustness: (i) label permutation within the in-context demonstration shots, and (ii) decision-rule sensitivity, contrasting the commonly reported HITS@3 with our primary metric, HITS@1. These tests probe robustness to spurious label-token mappings and to the choice of evaluation rule, respectively.
@@
Why permutation hurts 0.5B-1B but not 0.13B? At 0.13B, the robustness to label permutation suggests it benefits from demonstrations via format/topical priming and answer-frequency priors, but does not reliably exploit the labeltoken mapping. In contrast, 0.5B-1B models more strongly use the in-context mapping; permuting labels therefore contradicts a cue they have learned to trust, producing clear drops. This is consistent with reports that (i) labels in demos can be non-essential for smaller/less capable settings-format and priors often dominate (Min et al., 2022; Zhao et al., 2021), and (ii) the ability to override priors and follow contradictory, flipped labels emerges with scale (Wei et al., 2023).
@@
Following common practice for function-probe tasks in the Todd et al. (2024) suite-which reports HITS@3 (top-3 token accuracy)-we compare our primary decision rule (HITS@1) with HITS@3. ICL performance is summarized in Figure 7, and per-task HITS@3 accuracies are listed in Table 11. While HITS@3 increases absolute scores across the board, the relative ordering and gaps between variants remain effectively unchanged.